XYZ Co. ABC Co. $922,400 $922,400 518,600 313,100 Contribution margin 403,800 609,300 303,800 509,300 Operating income $100,000 $100,000 Sales revenue Variable costs Fixed costs Calculate each companys degree of operating leverage. (Round answers to 4 decimal places, e.g. 1.2531.) Degree of operating leverage 4.038 6.093 XYZ ABC Determine which companys cost structure makes it more sensitive to changes in its sales volume ABC LINK TO TEXT Determine the effect on each companys operating income if sales decrease by 10% XYZ: Operating income would decrease▼ by $ 40,380 to become $ 59,620 60,930 to become $ 39,070 ABC: Operating income would decrease by $
Answer #1

Degree of Operating Leverage = Contribution Margin/Operating Income

XYZ: 403,800/100,000 = 4.038

ABC: 609,300/100,000 = 6.093

A company with higher DOL has higher fixed costs and hence, operating income changes higher with changes in sales

Hence, company ABC is more sensitive to changes in sales volume

XYZ: Operating income would decrease by 4.038*100,000*10% = 40,380 to become 59,620

ABC: Operating Income would decrease by 60,930 to become 39,070